13.2x2+ nx + 5 = 0 has imaginary roots. The range of values of n is?


14.The range of values of, (x + 1)/ (x2 + 1)  X e R is?


1
Expert's answer
2021-02-28T16:59:49-0500

13. For any quadratic equation to have imaginary roots, b² – 4ac must be less than 0


b² – 4ac < 0

a = 2 ; b = n ; c = 5


n² – 4(2)(5) < 0

n² – 40 < 0

n² < 40

=> "-\sqrt{40} < n < \sqrt{40}"

= "-2\sqrt{10} < n < 2\sqrt{10}"


14. Let "y = \frac{x+1}{x²+1}"

Make x the subject

"yx\u00b2+y = x+1"

"yx\u00b2-x+y-1=0"

From the general quadratic formula

"ax\u00b2+bx+c=0"

We can say that

a = y ; b = -1; c = y-1


from the quadratic formula

"x=\frac{-b±\sqrt{b²-4ac}}{2a}"

"x=\frac{-(-1)±\sqrt{(-1)²-4(y)(y-1)}}{2y}"

"x=\frac{1±\sqrt{1-4(y²-y}}{2y}"

"x=\frac{1±\sqrt{1-4y²+4y}}{2y}"


For the above equation to be defined, then "1-4y\u00b2+4y\u22650"


-4y² + 4y + 1 ≥ 0

4y² - 4y - 1 ≤ 0

Divide through by 4

"y²-y-\frac{1}{4} ≤0"

"y²-y+\frac{1}{4}-\frac{1}{4}-\frac{1}{4}≤0"

"y²-y+\frac{1}{4}-\frac{1}{2}≤0"

Multiply through by 4

(4y² – 4y + 1) – 2 ≤ 0

(2y – 1)² – 2 ≤ 0

(2y – 1)² ≤ 2

=> "-\sqrt{2} ≤ 2y-1≤\sqrt{2}"

"-\sqrt{2}≤2y-1"

"2y-1≥-\sqrt{2}"

"y≥\frac{1-\sqrt{2}}{2}"


=>"2y-1≤\sqrt{2}"

"y ≤ \frac{\sqrt{2}+1}{2}"


The range of values is

"[\frac{1-\sqrt{2}}{2},0) \bigcup (0, \frac{\sqrt{2}+1}{2}]"